OpenStudy (anonymous):

Adding and subtracting radical expression, how do you solve 4√12 - √75

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you need to write both in simplest radical form

OpenStudy (mathstudent55):

Look inside each root, and factor out the largest perfect square you can.

OpenStudy (anonymous):

is it 4√6√2 - √25√3

OpenStudy (misty1212):

no

OpenStudy (misty1212):

the second one is close

OpenStudy (misty1212):

\[-5\sqrt3\] for the second one, since \(\sqrt{25}=5\)

OpenStudy (mathstudent55):

6 is not a perfect square, so it won't help. In the second root, 25 is good.

OpenStudy (anonymous):

Ooooh

OpenStudy (misty1212):

\[12=4\times 3\] so \[\sqrt{12}=\sqrt4\sqrt3=2\sqrt3\]

OpenStudy (misty1212):

the entire thing becomes \[8\sqrt3-5\sqrt3\]

OpenStudy (misty1212):

then combine like terms like \[8x-5x\]

OpenStudy (anonymous):

My goodness! It looks so complicated

OpenStudy (mathstudent55):